OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, -1).

OpenStudy (hoblos):

thats almost the same as the previous!! can you try to solve it?

OpenStudy (anonymous):

alright ill try my best lol

OpenStudy (hoblos):

tell us what you get so that we can check it ;)

OpenStudy (karatechopper):

parallel= same slope perpendicular= opposite of reciprocal Ex. slope= 1/4 perpendicular slope: -4/1

OpenStudy (anonymous):

ok so far i got it in slop int form : y=-3x+7 is that right?

OpenStudy (karatechopper):

yes

OpenStudy (anonymous):

i ended up with y = -3x - 3 correct?

OpenStudy (anonymous):

I see but wouldnt it be y = -3x + 17 ?

OpenStudy (hoblos):

let the line we want equation of be (d) and its slope is m slope of y=-3x+7 is m' = -3 m*m'=-1 so m = 1/3 y+1 = (1/3)(x-6) y+1 = (1/3)x -2 y = (1/3)x -1